OpenStudy (anonymous):

If a can of ravioli with a diameter of 2 inches and a height of 4.5 inches holds 12 ounces of ravioli, how many ounces of ravioli would a can with the same height, but double the diameter, hold?

OpenStudy (mertsj):

\[V _{1}=\pi(1)^2(4.5)=4.5\pi\] \[V _{2}=\pi(2)^2(4.5)=18\pi\]

OpenStudy (mertsj):

Well, look at that!! The bigger can has a volume that is 4 times the volume of the little can so I guess it would hold 4 times as much. What do you think?